OpenStudy (anonymous):

Solve -4(x + 10) - 6 = -3(x - 2). -40 -46 -52 52

OpenStudy (anonymous):

how u doit and explain

OpenStudy (anonymous):

-4(x+10)-6=-3(x-2) -4x-40-6=-3x+6 You want x on one side -40-6-6=-3x+4x -52=x x=-52

OpenStudy (anonymous):

thanks